# Flavored QCD axion and Modular invariance

### v2

@@ II.2 Gauged $U(1)_{X}$ anomaly-free and global $U(1)_{X}$
The effective action (29), after canonical normalization with $\theta_{X}=a_{\theta}/8\pi^{2}f_{\theta}$ where $f_{\theta}=\sqrt{2K_{U_{X}\bar{U}_{X}}}/8\pi^{2}$, and incorporating the gauge kinetic function, yields kinetic terms for both axions along with their couplings to the topological term ${\rm Tr}(Q^{\mu\nu}\tilde{Q}_{\mu\nu})$. The $U(1)_{X}$ gauge boson acquires mass $m_{X}=\sqrt{2K_{U_{X}\bar{U}_{X}}(\delta^{\rm GS}_{X}/16\pi^{2})^{2}+2f^{2}_{X}}$ through the super-Higgs mechanism, while the D-term potential $V_{D}$ remains. The open string axion $A_{X}$ (with its decay constant $f_{X}$) is mixed linearly with the closed string axion $a_{\theta}$ (with its decay constant $f_{\theta}$) such that the orthogonal combinations $G\approx a_{\theta}$ (NG mode, absorbed by the gauge boson) and $\tilde{A}\approx A_{X}$ (pseudo-NG mode, remaining as the physical axion) emerge when $f_{\theta}\gg f_{X}$. Below the scale $m_{X}$, the gauge boson decouples, leaving an anomalous  [^11] In certain gravitational backgrounds, non-perturbative quantum gravitational anomaly effects (specifically, instantons) can lead to an anomalous non-conservation of an axial vector current: $\partial_{\mu}J^{\mu}\propto R\tilde{R}$ where $R$ is the Riemann curvature tensor and $\tilde{R}$ is its dual Kamionkowski:1992mf . In this work, for simplicity, we assume that the gravitational mixed anomalies are canceled by the GS mechanism. global $U(1)_{X}$. See the details in Ref.Ahn:2023iqa ; Ahn:2016typ .
@@ II.3 The VEV of the modulus $\tau$
To determine Yukawa couplings, gauge couplings, SUSY-breaking scale, and cosmological constant, we consider a simple modular- and gauge-invariant superpotential $W(S,U_{X},\tau)$ in terms of the Dedekind $\eta$-function, which  [^13] Here, for simplicity, the Dedekind multiplier $e^{i\epsilon(a,b,c,d)}$ is omitted, where $\epsilon(a,b,c,d)$ is a moduli-independent phase, which can depend on the $SL(2,\mathbb{Z})$ transformation. Under the modular transformation given in Eq.(1), in general, the Dedekind eta function $\eta(\tau)$ transforms as $\eta(\tau)\rightarrow e^{i\epsilon}(c\tau+d)^{1/2}\eta(\tau)$ and the superpotential $W$ transforms as $W\rightarrow e^{-g(\tau)-i2h\epsilon}W$ where the matter fields $\varphi_{i}$ transform as: $\varphi_{i}\rightarrow e^{-i\epsilon_{i}}(c\tau+d)^{-k_{i}}\varphi_{i}$ with the condition $\sum_{i}\epsilon_{i}=\epsilon$, and the dilaton transforms as $S\rightarrow S-\frac{1}{4\pi^{2}}\{g(\tau)+i2h\epsilon\}$. is a modular form of weight $1/2$, $\eta(\tau)\rightarrow(c\tau+d)^{1/2}\eta(\tau)$, and non-perturbative effect for gaugino condensation:
$W(S,U_{X},\tau)=C_{0}\frac{e^{-\alpha S}M^{3}_{P}}{[\eta(\tau)]^{2h(1+\alpha/4\pi^{2})}}+\frac{M^{3}_{P}}{[\eta(\tau)]^{2h}}\big(Ae^{-aU_{X}}-Be^{-bU_{X}}\big)\,$, (33)
where $C_{0}$, $\alpha,a,b$ are constants. In the limit $\alpha\rightarrow 0$, this superpotential takes a racetrack type for $U_{X}$Kachru:2003aw . Under the modular transformation Eq.(1), the modular invariance of the generalized Kähler potential implies that the superpotential transforms as $W\rightarrow We^{-g(\tau)}$ with Eq.(6). Under the $U(1)_{X}$ transformation of $U_{X}$, see above Eq.(27), the $A(\varphi_{X}/M_{P})$ and $B(\varphi_{X}/M_{P})$, which are analytic function of $\varphi_{X}$, transform as
@@
$D_{\tau}W=-3W\Big[\frac{1-1/(4\pi^{2}y)}{\tau-\bar{\tau}}+2\frac{\eta^{\prime}(\tau)}{\eta(\tau)}\Big]-\frac{3}{2\pi^{2}}\frac{\eta^{\prime}(\tau)}{\eta(\tau)}\alpha C_{0}\frac{e^{-\alpha S}M^{3}_{P}}{[\eta(\tau)]^{6(1+\alpha/4\pi^{2})}}\,$,
$D_{S}W=-\frac{W}{y}-\alpha C_{0}\frac{e^{-\alpha S}M^{3}_{P}}{[\eta(\tau)]^{6(1+\alpha/4\pi^{2})}}\,$, (35)
where $y=S+\bar{S}-\frac{3}{4\pi^{2}}\ln(-i\tau+i\bar{\tau})$. For small $\alpha$, the superpotential Eq.(33) becomes
$W$ $=$ $W(U_{X},\tau)-\alpha\,C_{0}\frac{M^{3}_{P}}{[\eta(\tau)]^{6}}\big(S+\frac{3}{2\pi^{2}}\ln\eta(\tau)\big)$ (36)
$+$ $\frac{1}{2}\alpha^{2}C_{0}\frac{M^{3}_{P}}{[\eta(\tau)]^{6}}\big(S+\frac{3}{2\pi^{2}}\ln\eta(\tau)\big)^{2}-\frac{1}{6}\alpha^{3}C_{0}\frac{M^{3}_{P}}{[\eta(\tau)]^{6}}\big(S+\frac{3}{2\pi^{2}}\ln\eta(\tau)\big)^{3}+.$...
@@
where $A_{0}$ and $B_{0}$ are the values of $A(\varphi_{X}/M_{P})$ and $B(\varphi_{X}/M_{P})$ at $\langle\varphi_{X}\rangle$, respectively.
@@
$V^{(2)}_{F}$ $=$ $e^{K/M^{2}_{P}}\frac{M^{6}_{P}}{|\eta(\tau)|^{12}}|C_{0}|^{2}\Big\{|S+\frac{3}{2\pi^{2}}\ln\eta(\tau)|^{2}\big(K^{\tau\bar{\tau}}|H|^{2}+\frac{1}{M^{2}_{P}}\big)$ (42)
- $\frac{y}{M^{2}_{P}}\big(S+\bar{S}+\frac{3}{2\pi^{2}}\ln\eta(\tau)\eta(\bar{\tau})\big)+\frac{y^{2}}{M^{2}_{P}}$
- $\frac{3}{2\pi^{2}}K^{\tau\bar{\tau}}\big((S+\frac{3}{2\pi^{2}}\ln\eta(\tau))H\frac{\eta^{\prime}(\bar{\tau})}{\eta(\bar{\tau})}+h.c.\big)+\frac{2}{\pi^{4}}K^{\tau\bar{\tau}}\big|\frac{\eta^{\prime}(\tau)}{\eta(\tau)}\big|^{2}\Big\}\,$,
@@
$\Big|\frac{F^{\tau}}{F^{S}}\Big|\approx\frac{3}{2\pi^{2}}\frac{\langle K^{\tau\bar{\tau}}\rangle}{\langle K^{S\bar{S}}\rangle}\Big|\frac{\eta^{\prime}(\tau_{0})}{\eta(\tau_{0})}\Big|\Big|\frac{2s_{0}-\frac{3}{4\pi^{2}}\ln(2{\rm Im}\,\tau_{0})}{-s_{0}+\frac{3}{4\pi^{2}}\big(2\ln\eta(\tau_{0})+\ln(2{\rm Im}\,\tau_{0})\big)}\Big|\ll 1\,$, (44)
where $K^{S\bar{S}}=y^{2}/M^{2}_{P}$, $K^{\tau\bar{\tau}}=-(\tau-\bar{\tau})^{2}/\{6M^{2}_{P}\big(1-1/(4\pi^{2}y)+3/(8\pi^{4}y^{2})\big)\}$, and $|\eta^{\prime}(i)|\approx 0.192$. This implies that supersymmetry is broken predominantly by the dilaton $S$ and slightly by the modulus $\tau$, induced by $e^{-\alpha S}$ term in Eq.(33). This is decoupled from the AdS minimum of the $U_{X}$ direction. When including the uplifting term and $\alpha\neq 0$, the gravitino mass $m^{2}_{3/2}=e^{K/M^{2}_{P}}|W|^{2}/M^{4}_{P}$ becomes
@@ III Superpotential for Quark and Lepton
[…]
@@
assuming $\langle\chi\rangle,\langle\tilde{\chi}\rangle\gg\langle H_{u,d}\rangle$. This supersymmetric solution satisfies the $D$-flatness condition for $\xi^{\rm FI}_{X}=0$. For $\xi^{\rm FI}_{X}\neq 0$, SUSY must be broken by the $F$-term to allow the $D$-term to uplift the potential from the AdS minimum to dS Achucarro:2006zf ; Burgess:2003ic (see Sec.II.3).
@@ III.1 Modular-invariant Yukawa superpotential for quark
where all Yukawa coefficients $\alpha_{i}$ are complex numbers with unit magnitude, and dots represent higher-order contributions compactly expressed as $\sum^{\infty}_{n=1}(\frac{\chi\tilde{\chi}}{\Lambda^{2}})^{n}\times\text{\it leading terms}$. Here the flavor dynamics scale $\Lambda$ can be identified with the stabilization scale of the lightest modulus $m_{\rm modulus}\sim\Lambda$, ensuring that fluctuations of heavy moduli are suppressed below $\Lambda$. These corrections modify the effective Yukawa coefficients $\alpha_{i}$, constrained by
@@
${\cal L}_{\vartheta}=\Big(\vartheta_{\rm eff}+\frac{A_{X}}{F_{a}}\Big)\frac{\alpha_{s}^{\prime}}{8\pi}G^{a\mu\nu}\tilde{G}^{a}_{\mu\nu}\qquad\text{with}~F_{a}=\frac{f_{A}}{\delta^{G}_{X}}\,$, (53)
@@ IV Quark, Lepton, and Flavored-QCD axion
At energies below the electroweak scale when $H_{u(d)}$ acquire non-zero VEVs all quarks and leptons obtain masses. The relevant quark and lepton interactions are given from Eqs.(50,55,56) by
@@ IV.1 Quark and charged-lepton masses, mixing, and QCD axion interactions
After diagonalizing the quark mass matrices of Eqs.(64,68), the leading-order flavored-QCD axion interactions with quarks (up to $\mathcal{O}(\lambda^{4})$) are  [^16] The left-handed quark and the right-handed charged-lepton mixing matrices, $V^{d(u)}_{L}$ and $V^{\ell}_{R}$, are determined in Appendix E. These matrices govern the flavored-QCD axion couplings to down-type quarks, up-type quarks, and charged leptons, respectively.
@@ IV.2 Neutrino mass and mixing
The seesaw mechanism in Eq.(96) operates at the $U(1)_{X}$ symmetry breakdown scale, while its implications are measured by experiments below the electroweak scale. Therefore, quantum corrections to neutrino masses and mixing angles can be crucial, especially for degenerate neutrino masses Antusch:2005gp . However, based on our numerical calculation that the neutrino mass spectra exhibit normal mass hierarchy (NH) at the scale of $U(1)_{X}$ breakdown (as depicted in the left panel of Fig.3), we can safely assume that the renormalization group running effect on observables can be ignored. To show the viable parameter space, we scan the precision constraints $\{\theta_{13}$, $\theta_{23}$, $\theta_{12}$, $\Delta m^{2}_{\rm Sol}$, $\Delta m^{2}_{\rm Atm}\}$ at $3\sigma$ from Table-3. Using the reference values from Eqs.(41,94) and taking $\langle\chi\rangle=2\times 10^{10}$ GeV of Eq.(92), see also Eq.(97), we determine the input parameter spaces in Eqs.(77) and (81) at the $U(1)_{X}$ breaking scale :

### v3

@@ II.2 Gauged $U(1)_{X}$ anomaly-free and global $U(1)_{X}$
The effective action (29), after canonical normalization with $\theta_{X}=a_{\theta}/8\pi^{2}f_{\theta}$ where $f_{\theta}=\sqrt{2K_{U_{X}\bar{U}_{X}}}/8\pi^{2}$, and incorporating the gauge kinetic function, yields kinetic terms for both axions along with their couplings to the topological term ${\rm Tr}(Q^{\mu\nu}\tilde{Q}_{\mu\nu})$. The $U(1)_{X}$ gauge boson acquires mass $m_{X}=\tilde{g}_{X}\sqrt{2K_{U_{X}\bar{U}_{X}}(\delta^{\rm GS}_{X}/16\pi^{2})^{2}+2f^{2}_{X}}$ through the super-Higgs mechanism, while the D-term potential $V_{D}$ remains. The open string axion $A_{X}$ (with its decay constant $f_{X}$) is mixed linearly with the closed string axion $a_{\theta}$ (with its decay constant $f_{\theta}$) such that the orthogonal combinations $G\approx a_{\theta}$ (NG mode, absorbed by the gauge boson) and $\tilde{A}\approx A_{X}$ (pseudo-NG mode, remaining as the physical axion) emerge when $f_{\theta}\gg f_{X}$. Below the scale $m_{X}$, the gauge boson decouples, leaving an anomalous  [^10] In certain gravitational backgrounds, non-perturbative quantum gravitational anomaly effects (specifically, instantons) can lead to an anomalous non-conservation of an axial vector current: $\partial_{\mu}J^{\mu}\propto R\tilde{R}$ where $R$ is the Riemann curvature tensor and $\tilde{R}$ is its dual Kamionkowski:1992mf . In this work, for simplicity, we assume that the gravitational mixed anomalies are canceled by the GS mechanism. global $U(1)_{X}$. See the details in Ref.Ahn:2023iqa ; Ahn:2016typ .
@@ II.3 The VEV of the modulus $\tau$
To determine Yukawa couplings, gauge couplings, SUSY-breaking scale, and cosmological constant, we consider a simple modular- and gauge-invariant superpotential $W(S,U_{X},\tau)$ in terms of the Dedekind $\eta$-function, which  [^12] Here, for simplicity, the Dedekind multiplier $e^{i\epsilon(a,b,c,d)}$ is omitted, where $\epsilon(a,b,c,d)$ is a moduli-independent phase, which can depend on the $SL(2,\mathbb{Z})$ transformation. Under the modular transformation given in Eq.(1), in general, the Dedekind eta function $\eta(\tau)$ transforms as $\eta(\tau)\rightarrow e^{i\epsilon}(c\tau+d)^{1/2}\eta(\tau)$ and the superpotential $W$ transforms as $W\rightarrow e^{-g(\tau)-i2h\epsilon}W$ where the matter fields $\varphi_{i}$ transform as: $\varphi_{i}\rightarrow e^{-i\epsilon_{i}}(c\tau+d)^{-k_{i}}\varphi_{i}$ with the condition $\sum_{i}\epsilon_{i}=\epsilon$, and the dilaton transforms as $S\rightarrow S-\frac{1}{16\pi^{2}}\{g(\tau)+i2h\epsilon\}$. is a modular form of weight $1/2$, $\eta(\tau)\rightarrow(c\tau+d)^{1/2}\eta(\tau)$, and non-perturbative effect for gaugino condensation:
$W(S,U_{X},\tau)=C_{0}\frac{e^{-\alpha S}M^{3}_{P}}{[\eta(\tau)]^{2h(1+\alpha/16\pi^{2})}}+\frac{M^{3}_{P}}{[\eta(\tau)]^{2h}}\big(Ae^{-aU_{X}}-Be^{-bU_{X}}\big)\,$, (34)
where $C_{0}$, $\alpha,a,b$ are constants. Here  [^13] The gauge kinetic function in Eq.(4) applies to the visible sector. the non-perturbative contributions originate from two distinct hidden gauge sectors: one condensing hidden gauge group with gauge kinetic function $f_{1}=S$, generating the first exponential term in $S$, and another hidden gauge group with gauge kinetic function $f_{2}=U_{X}$, generating the second exponential term in $U_{X}$. In the limit $\alpha\rightarrow 0$, this superpotential takes a racetrack type for $U_{X}$Kachru:2003aw . Under the modular transformation Eq.(1), the modular invariance of the generalized Kähler potential implies that the superpotential transforms as $W\rightarrow We^{-g(\tau)}$ with Eq.(6). Under the $U(1)_{X}$ transformation of $U_{X}$, see above Eq.(27), the $A(\varphi_{X}/M_{P})$ and $B(\varphi_{X}/M_{P})$, which are analytic function of $\varphi_{X}$, transform as
@@
$D_{\tau}W=-3W\Big[\frac{1-1/(16\pi^{2}y)}{\tau-\bar{\tau}}+2\frac{\eta^{\prime}(\tau)}{\eta(\tau)}\Big]-\frac{3}{8\pi^{2}}\frac{\eta^{\prime}(\tau)}{\eta(\tau)}\alpha C_{0}\frac{e^{-\alpha S}M^{3}_{P}}{[\eta(\tau)]^{6(1+\alpha/16\pi^{2})}}\,$,
$D_{S}W=-\frac{W}{y}-\alpha C_{0}\frac{e^{-\alpha S}M^{3}_{P}}{[\eta(\tau)]^{6(1+\alpha/16\pi^{2})}}\,$, (36)
where $y=S+\bar{S}-\frac{3}{16\pi^{2}}\ln(-i\tau+i\bar{\tau})$. For small $\alpha$, the superpotential Eq.(34) becomes
$W$ $=$ $W(U_{X},\tau)-\alpha\,C_{0}\frac{M^{3}_{P}}{[\eta(\tau)]^{6}}\big(S+\frac{3}{8\pi^{2}}\ln\eta(\tau)\big)$ (37)
$+$ $\frac{1}{2}\alpha^{2}C_{0}\frac{M^{3}_{P}}{[\eta(\tau)]^{6}}\big(S+\frac{3}{8\pi^{2}}\ln\eta(\tau)\big)^{2}-\frac{1}{6}\alpha^{3}C_{0}\frac{M^{3}_{P}}{[\eta(\tau)]^{6}}\big(S+\frac{3}{8\pi^{2}}\ln\eta(\tau)\big)^{3}+.$...
@@
where $A_{0}$ and $B_{0}$ are the values of $A(\varphi_{X}/M_{P})$ and $B(\varphi_{X}/M_{P})$ at $\langle\varphi_{X}\rangle$, respectively. The $U(1)_{X}$ charged fields $\varphi_{X}$ are given by $\chi$ and $\tilde{\chi}$ introduced in Sec.III. As shown in Appendix C, the flavon fields $\varphi_{X}=\{\chi,\tilde{\chi}\}$ are stabilized by the leading-order conditions $D_{\chi}W=D_{\tilde{\chi}}W=D_{\chi_{0}}W=0$ together with the $U(1)_{X}$ D-flatness. The resulting VEVs in Eq.(114) determine constant coefficients $A_{0}$ and $B_{0}$.
@@
$V^{(2)}_{F}$ $=$ $e^{K/M^{2}_{P}}\frac{M^{6}_{P}}{|\eta(\tau)|^{12}}|C_{0}|^{2}\Big\{|S+\frac{3}{8\pi^{2}}\ln\eta(\tau)|^{2}\big(K^{\tau\bar{\tau}}|H|^{2}+\frac{1}{M^{2}_{P}}\big)$ (43)
- $\frac{y}{M^{2}_{P}}\big(S+\bar{S}+\frac{3}{8\pi^{2}}\ln\eta(\tau)\eta(\bar{\tau})\big)+\frac{y^{2}}{M^{2}_{P}}$
- $\frac{3}{8\pi^{2}}K^{\tau\bar{\tau}}\big((S+\frac{3}{8\pi^{2}}\ln\eta(\tau))H\frac{\eta^{\prime}(\bar{\tau})}{\eta(\bar{\tau})}+h.c.\big)+\frac{9}{64\pi^{4}}K^{\tau\bar{\tau}}\big|\frac{\eta^{\prime}(\tau)}{\eta(\tau)}\big|^{2}\Big\}\,$,
@@
$\Big|\frac{F^{\tau}}{F^{S}}\Big|\approx\frac{3}{8\pi^{2}}\frac{\langle K^{\tau\bar{\tau}}\rangle}{\langle K^{S\bar{S}}\rangle}\Big|\frac{\eta^{\prime}(\tau_{0})}{\eta(\tau_{0})}\Big|\Big|\frac{2s_{0}-\frac{3}{16\pi^{2}}\ln(2{\rm Im}\,\tau_{0})}{-s_{0}+\frac{3}{16\pi^{2}}\big(2\ln\eta(\tau_{0})+\ln(2{\rm Im}\,\tau_{0})\big)}\Big|\ll 1\,$, (45)
where $K^{S\bar{S}}=y^{2}/M^{2}_{P}$, $K^{\tau\bar{\tau}}=-(\tau-\bar{\tau})^{2}/\{6M^{2}_{P}\big(1-1/(16\pi^{2}y)+3/(128\pi^{4}y^{2})\big)\}$, and $|\eta^{\prime}(i)|\approx 0.192$. This implies that supersymmetry is broken predominantly by the dilaton $S$ and slightly by the modulus $\tau$, induced by $e^{-\alpha S}$ term in Eq.(34). This is decoupled from the AdS minimum of the $U_{X}$ direction. When including the uplifting term and $\alpha\neq 0$, the gravitino mass $m^{2}_{3/2}=e^{K/M^{2}_{P}}|W|^{2}/M^{4}_{P}$ becomes
@@ III Superpotential for Quark and Lepton
where $g_{\chi_{0}}$ and $g_{\chi}$ are initially set to unity but receive corrections from higher-order terms (see Eq.(52)). The scale $\mu_{\chi}$ sets the spontaneous $U(1)_{X}$ breaking scale. Minimizing the $F$-term potential (Appendix C) yields
@@
assuming $\langle\chi\rangle,\langle\tilde{\chi}\rangle\gg\langle H_{u,d}\rangle$. This supersymmetric solution automatically satisfies the $D$-flatness condition for $\xi^{\rm FI}_{X}=0$. When $\xi^{\rm FI}_{X}\neq 0$ below the string scale, SUSY must be broken via the $F$-terms to allow the $D$-term to act as an uplift, raising the vacuum from the AdS minimum to dS vacuum Achucarro:2006zf ; Burgess:2003ic (see Sec.II.3).
@@ III.1 Modular-invariant Yukawa superpotential for quark
where all Yukawa coefficients $\alpha_{i}$ are complex numbers with unit magnitude, and dots represent higher-order contributions compactly expressed as $\sum^{\infty}_{n=1}(\frac{\chi\tilde{\chi}}{\Lambda^{2}})^{n}\times\text{\it leading terms}$. Here the flavor dynamics scale $\Lambda$ can be identified with the string scale, where the bare gauge coupling at $\Lambda$ is given by $1/g^{2}_{Q}(\Lambda)=\kappa_{Q}/g^{2}_{s}+\tilde{\kappa}_{Q}/g^{2}_{X}$, see Eq.(4) and above Eq.(30). These corrections modify the effective Yukawa coefficients $\alpha_{i}$, constrained by
@@
${\cal L}_{\vartheta}=\Big(\vartheta_{\rm eff}+\frac{A_{X}}{F_{a}}\Big)\frac{\alpha_{s}^{\prime}}{16\pi}G^{a\mu\nu}\tilde{G}^{a}_{\mu\nu}\qquad\text{with}~F_{a}=\frac{f_{A}}{\delta^{G}_{X}}\,$, (54)
@@ IV Quark, Lepton, and Flavored-QCD axion
At energies below the electroweak scale when $H_{u(d)}$ acquire non-zero VEVs all quarks and leptons obtain masses. The relevant quark and lepton interactions are given from Eqs.(51,56,57) by
@@ IV.1 Quark and charged-lepton masses, mixing, and QCD axion interactions
After diagonalizing the quark mass matrices of Eqs.(65,69), the leading-order flavored-QCD axion interactions with quarks (up to $\mathcal{O}(\lambda^{4})$) are  [^15] The left-handed quark and the right-handed charged-lepton mixing matrices, $V^{d(u)}_{L}$ and $V^{\ell}_{R}$, are determined in Appendix E. These matrices govern the flavored-QCD axion couplings to down-type quarks, up-type quarks, and charged leptons, respectively.
@@ IV.2 Neutrino mass and mixing
The seesaw mechanism in Eq.(97) operates at the $U(1)_{X}$ symmetry breakdown scale, while its implications are measured by experiments below the electroweak scale. Therefore, quantum corrections to neutrino masses and mixing angles can be crucial, especially for degenerate neutrino masses Antusch:2005gp . However, based on our numerical calculation that the neutrino mass spectra exhibit normal mass hierarchy (NH) at the scale of $U(1)_{X}$ breakdown (as depicted in the left panel of Fig.3), we can safely assume that the renormalization group running effect on observables can be ignored. To show the viable parameter space, we scan the precision constraints $\{\theta_{13}$, $\theta_{23}$, $\theta_{12}$, $\Delta m^{2}_{\rm Sol}$, $\Delta m^{2}_{\rm Atm}\}$ at $3\sigma$ from Table-3. Using the reference values from Eqs.(42, 95) and taking $\langle\chi\rangle=2\times 10^{10}$ GeV of Eq.(93), see also Eq.(98), we determine the input parameter spaces in Eqs.(78) and (82) at the $U(1)_{X}$ breaking scale :